Operations Report – Feb 28th

Crew 222 Operations Report 28FEB2020

SOL: 12

Name of person filing report: Valentin PRUDHOMME

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Thursday generator:

Off at 5:30am

On at 9:00pm SOC was 83%

Friday generator:

Off at 5:00am

On at 9:00pm SOC was XX%

Diesel Reading – 70%

Station Propane Reading – 83%

Ethanol Free Gasoline – 7 gallons

Water (loft tank): 47 gallons

Water Meter: 149590.24 units

Water (static tank) – 176 gallons

Static to Loft Pump used – yes

Water in GreenHab – 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: yes

Deimos rover used: no

Hours: 146.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Nothing to report

Beginning charge: Nothing to report

Ending charge: Nothing to report

Currently charging: Nothing to report

Spirit rover: no

Hours: 132.8

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes.

Hours: Nothing to report

Beginning charge: Nothing to report

Ending charge: Nothing to report

Currently charging: Nothing to report

Curiosity rover used: no

Hours: 138.8

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes (Honda, 1, 2 and 3)

Reason for use: Run during the engineer EVA for the battery and for the

EVA this morning.

Oil Added? no

ATV Fuel Used: 20% of the tank.

Hours the ATVs were Used today: 45 minutes

Notes on ATVs: Oil checked. The Honda and the ATV number 1’s oil light lit up. At the end of the EVA, the red one couldn’t change gear. I stopped and it was stuck in 5th gear. We had to push it to change gear. When we finally put it in neutral, we started it again and it worked well until the end (the gears were working fine).

HabCar used and why, where? The HabCar is now in the shop awaiting repairs.

CrewCar used and why, where? The CrewCar is in Grand Junction.

General notes and comments: Nothing to report

Summary of internet: All ok.

Summary of suits and radios: Suits 7, 8, 9, 10 were used during the EVA. The suit number 7 was working well at the beginning but stopped working in the middle of the EVA. We fixed this in the RAM during this afternoon and now it’s working well.

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Fixing of the suit number 7. It’s working well now. Finished the spacesuit experiment and all working well now.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Journalist Report – Feb 28th

Crew 222 Journalist Report 28Feb2020

Author: Marie Bochard, Crew Journalist

Sol 12

Title: First breath of fresh air

To enjoy fully our last simulation day, we started the day off with a long EVA. Nina, Eishi, Valentin and I each took an ATV and headed North to the Moons and Copernicus Highway (a dust path, as the name doesn’t suggest). The landscape was incredible. Every step we took, the ground was a different color. It really felt like we were exploring an alien planet. We found plenty of fossils of oysters, proof that there was plenty of water here a long time ago. There were glass-like shards on the floor, possibly made by lightning hitting the sand… that’s our guess. We surely walked where no one has every walked before, and it

created a very unique mix between feeling special and feeling small. After a couple hours of exploring, we headed back on the ATVs and everyone enjoyed the ride to the Hab.

The rest of the day was mostly spent cleaning. We collected our belongings and started packing.

Then it was over. We opened the Hab door at 5:30pm, finally breathing fresh air after two weeks here. How lucky are we Earthlings to be able to breathe outside and to walk freely without having to wear a pressurized suit!

This mission has gone by tremendously fast. It feels like we just got here, and yet it’s been ages since we thought of the world outside. Everyone enjoyed the stay here, and we feel closer than ever. Missions like this one can create ever-lasting bonds, sometimes more than years can achieve. It will be forged in everyone’s memory for the rest of our lives, I can guarantee that.

Operations Report – Feb 27th

Crew 222 Operations Report 27FEB2020

SOL: 11

Name of person filing report: Valentin PRUDHOMME

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Wednesday generator:

Off at 6:00am

On at 9:00pm SOC was 85%

Thursday generator:

Off at 5:30am

On at 9:00pm SOC was 83%

Diesel Reading – 72%

Station Propane Reading – 84%

Ethanol Free Gasoline – 7 gallons

Water (loft tank): 200 gallons

Water Meter: 149548.32 units

Water (static tank) – 218 gallons

Static to Loft Pump used – yes

Water in GreenHab – 35.5 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: yes

Hours: 146.7

Beginning charge: 100%

Ending charge: 89%

Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Nothing to report

Beginning charge: Nothing to report

Ending charge: Nothing to report

Currently charging: Nothing to report

Spirit rover: no

Hours: 132.8

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes.

Hours: Nothing to report

Beginning charge: Nothing to report

Ending charge: Nothing to report

Currently charging: Nothing to report

Curiosity rover used: no

Hours: 138.8

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes (Honda, 1 and 2)

Reason for use: Run during the engineer EVA for the battery and for the

EVA this morning.

Oil Added? no

ATV Fuel Used: 20% of the tank.

Hours the ATVs were Used today: 30 minutes

Notes on ATVs: Oil checked. The Honda needs an oil change. The four ATVs

were operational this afternon.

HabCar used and why, where? The habcar is now in the shop awaiting repairs.

CrewCar used and why, where? The CrewCar is in Grand Junction.

General notes and comments: Nothing to report

Summary of internet: All ok.

Summary of suits and radios: Suits 7, 8, 9, 10 were used during the EVA.

The suit number 7 was working well at the beginning but two hours after the beginning of the EVA the fan was stopping at random moments but was restarting after. The fan never stopped totally but wasn’t working optimally.

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

GreenHab Report – Feb 28th

Crew 222 GreenHab Report – 28FEB2020

GreenHab Officer : Bleuenn ROIRON

Environmental control : Both heating and ambient, door open during the afternoon

Average temperatures: 21°C

Hours of supplemental light: 19h30 PM to 00h

Daily water usage for crops: 10,6 ( 6 from the static tank since the Blue tank is empty)

Daily water usage for research and/or other purposes: 0 Liters

Water in blue Tank: 0 gallons

Time(s) of watering crops: 2

Changes to crops: NA

Narrative:This morning, I went to the Greenhab to water plants, I uses water of the blue tank until it was empty so I started using water from the static tank.

Tonight i watered again plants, I also harvested some herbs for our tonight pie.

Harvest:

Oregano: 10g

Chives:8g

Sage:2g

Basil:13g

Raddishes:7g

Cilantro:5g

Support/supplies needed: There is no water left in the blue tank and I’ll have to use water from the static tank again tomorrow.

EVA Report – Feb 28th

Crew 222 EVA Report 28FEB2020

EVA # 13

Author: Marie BOCHARD (Journalist)

Purpose of EVA: Change batteries of LOAC and MegaARES, and then exploration of the Copernicus area

Start time: 9:15 AM

End time: 12:20 PM

Narrative: First, we checked all the rovers and the water level. We then took the ATVs to the experiments site and changed the batteries for the LOAC and the MegaARES in teams of two. Once that was all done, two of us we went back to the Hab with the batteries and left them in front of the Ram, because carrying them by ATV would have been dangerous. Once they were back, we headed up North to the Moons and Copernicus. We rode the ATVs on Brahe Hwy, but stopped soon after because the terrain was too rough for me to feel comfortable using an ATV. We stayed in Copernicus about 2 hours, exploring the area and enjoying our last EVA. Suit n°7 stopped working in the middle of the EVA, but this problem was solved by our Engineer once we were back (battery unplugged). At about 11:50 AM, we were back at the ATVs and headed back to the Hab. The Honda ATV had trouble restarting but worked in the end. We entered the station at 12:20 PM.

Destination: Copernicus Hwy

Coordinates (use UTM NAD27 CONUS): 4254500-517000

Participants: Marie BOCHARD (Journalist, EVA leader), Nina SEBDON

(Scientist), Eishi KIM (AST), Valentin PRUDHOMME (Crew engineer)

Road(s) and routes per MDRS Map: Cow Dung Road, Brahe Hwy

Mode of travel: ATVs: 1, 2, 3, Honda

Sol Summary Report – Feb 28th

Crew 222 Sol Summary Report 28FEB2020

Sol: 12

Summary Title: Last martian day

Author’s name: Benjamin Auzou

Mission Status: Sim ended, ready to receive crew 223

Sol Activity Summary: In the morning, the crew woke up with the last matinal workout session. Then four of us went on the last EVA, exploring the area of Copernicus. The day was very similar to our other sols, but it was the last !

In the afternoon, we cleaning the Hab in order to prepare the crew transition. At 6pm, we opened the airlock that we closed at the same hour two weeks ago. We all took a deep breath, we are back on Earth after two wonderful weeks on Mars that taught us a lot.

Look Ahead Plan:

Tomorrow, we will wake up very early to enjoy the sunrise from the ridge. Then we will prepare the arrival of crew 223, prepare them a nice martian lunch, get water and food for their rotation. This day will be intense.

Anomalies in work: none

Weather: sunny

Crew Physical Status: all members safe and sound

EVA: Exploration of Hab Ridge

Reports to be filed: Journalist report, Operations report, GreenHab Report, Astronomy Report, EVA Report, Photos Report, Photo of the day

Support Requested: nothing to report

Astronomy Report – Feb 28th

Crew 222 Astronomy Report 28 Feb 2020

Crew Astronomer: Eishi Kim

MDRS ROBOTIC OBSERVATORY

Robotic Telescope Requested (choose one) MDRS-14

Objects to be Imaged this Evening: Back to the research project, the
following targets are in the queue despite a cloudy sky: NGC 1961, NGC
925, NGC 4490, NGC 3368, NGC 3115, NGC 598, IC 342, NGC 5248, NGC 4038,
NGC 6503, NGC 2403.

Images submitted with this report: None

Problems Encountered: None

MUSK OBSERVATORY

Solar Features Observed: N/A

Images submitted with this report: N/A

Problems Encountered: N/A